# Insights into the Lottery Ticket Hypothesis and the Iterative Magnitude Pruning

### v1

@@ IV Results and Findings
$L$
Solution
$\lambda_{max}$
$V^{{}^{\prime}}(100)$
1
$W^{Pr{(min\_(0))}}_{(1)}$
1416.091
409.242
$W_{(1)}^{(min\_(1))}$
1590.829
407.417
2
$W^{Pr{(min\_(1))}}_{(2)}$
1512.426
402.678
$W_{(2)}^{(min\_(2))}$
1349.951
399.680
3
$W^{Pr{(min\_(2))}}_{(3)}$
1401.285
400.855
$W_{(3)}^{(min\_(3))}$
1194.164
386.948
4
$W^{Pr{(min\_(3))}}_{(4)}$
1266.607
392.829
$W_{(4)}^{(min\_(4))}$
1129.884
375.590
5
$W^{Pr{(min\_(4))}}_{(5)}$
17779.826
417.984
$W_{(5)}^{(min\_(5))}$
1094.078
372.753
6
$W^{Pr{(min\_(5))}}_{(6)}$
1590.785
420.431
$W_{(6)}^{(min\_(6))}$
1065.735
373.472
7
$W^{Pr{(min\_(6))}}_{(7)}$
2368.243
446.758
$W_{(7)}^{(min\_(7))}$
1065.881
375.401
8
$W^{Pr{(min\_(7))}}_{(8)}$
1882.278
432.505
$W_{(8)}^{(min\_(8))}$
1200.277
382.051
9
$W^{Pr{(min\_(8))}}_{(9)}$
3753.991
488.004
$W_{(9)}^{(min\_(9))}$
1334.666
394.166
10
$W^{Pr{(min\_(9))}}_{(10)}$
5426.848
525.621
$W_{(10)}^{(min\_(10))}$
1575.910
408.810
@@
Solution
Average Radius of Basin
$W_{(10)}^{(min\_(10))}$
0.232314
$W^{Pr{(min\_(9))}}_{(10)}$
0.183428
@@
$L$
Solution
$\lambda_{max}$
$V^{{}^{\prime}}(100)$
1
$W^{RPr{(min\_(1))}}_{(0)}$
1742.268
418.034
$W_{(0)}^{(min\_(0))}$
1575.277
415.42
2
$W^{RPr{(min\_(2))}}_{(1)}$
1572.214
415.826
$W_{(1)}^{(min\_(1))}$
1590.829
407.417
3
$W^{RPr{(min\_(3))}}_{(2)}$
1496.211
403.674
$W_{(2)}^{(min\_(2))}$
1349.951
399.68
4
$W^{RPr{(min\_(4))}}_{(3)}$
1326.676
394.347
$W_{(3)}^{(min\_(3))}$
1194.164
386.948
5
$W^{RPr{(min\_(5))}}_{(4)}$
1291.157
391.093
$W_{(4)}^{(min\_(4))}$
1129.884
375.59
6
$W^{RPr{(min\_(6))}}_{(5)}$
1316.275
391.471
$W_{(5)}^{(min\_(5))}$
1094.078
372.753
7
$W^{RPr{(min\_(7))}}_{(6)}$
1418.483
394.086
$W_{(6)}^{(min\_(6))}$
1065.735
373.472
8
$W^{RPr{(min\_(8))}}_{(7)}$
1464.987
398.559
$W_{(7)}^{(min\_(7))}$
1065.881
375.401
9
$W^{RPr{(min\_(9))}}_{(8)}$
1619.784
411.997
$W_{(8)}^{(min\_(8))}$
1200.277
382.051
10
$W^{RPr{(min\_(10))}}_{(9)}$
1913.214
427.475
$W_{(9)}^{(min\_(9))}$
1334.66
394.166
@@
Solution
Average Radius of Basin
$W_{(9)}^{(min\_(9))}$
0.091809
$W^{RPr{(min\_(10))}}_{(9)}$
0.074714
@@ IV-D Result 4: Why does the initialization proposed by the lottery ticket hypothesis work well?
Solution
$\lambda_{max}$
$V^{{}^{\prime}}(100)$
$W^{(RIPN)}$
6271.873
497.099
$W^{(min\_(10))}_{(10)}$
1575.910
408.810
@@ IV-E Result 5: Why do we need the Iterative Process, and why does one-shot pruning not work as well?
Solution
$\lambda_{max}$
$V^{{}^{\prime}}(100)$
$W^{(one\_shot)}$
2944.881
450.409
$W^{(min\_(10))}_{(10)}$
1575.910
408.810
@@ IV-F Result 6: What happens when you prune the smaller weights?
Solution
$\lambda_{max}$
$V^{{}^{\prime}}(100)$
$W^{(RPN\_2)}$
5635.883
487.644
$W^{(one\_shot)}$
2944.881
450.409
@@
Solution
$\lambda_{max}$
$V^{{}^{\prime}}(100)$
$W^{(RPN\_1)}$
2100.56
435.624
$W^{(min\_(10))}_{(10)}$
1575.91
408.81
@@ V Conclusion and Scope for Further Work
$L$
Solution
$\lambda_{max}$
$V^{{}^{\prime}}(100)$
1
$W^{Pr{(min\_(0))}}_{(1)}$
70.099
212.779
$W_{(1)}^{(min\_(1))}$
50.77
203.681
6
$W^{Pr{(min\_(5))}}_{(6)}$
86.27
198.106
$W_{(6)}^{(min\_(6))}$
37.08
177.233
11
$W^{Pr{(min\_(10))}}_{(11)}$
1413.523
414.844
$W_{(11)}^{(min\_(11))}$
1326.092
352.86
@@
$L$
Solution
$\lambda_{max}$
$V^{{}^{\prime}}(100)$
1
$W^{RPr{(min\_(1))}}_{(0)}$
43.777
219.996
$W_{(0)}^{(min\_(0))}$
67.929
220.915
6
$W^{RPr{(min\_(6))}}_{(5)}$
46.733
201.835
$W_{(5)}^{(min\_(5))}$
77.396
189.044
12
$W^{RPr{(min\_(12))}}_{(11)}$
4379.201
426.294
$W_{(11)}^{(min\_(11))}$
1326.092
352.86
@@
Solution
$\lambda_{max}$
$V^{{}^{\prime}}(100)$
$W^{(RPN)}$
51190.094
451.384
$W^{(min\_(12))}_{(12)}$
3992.4783
389.957

### v2

# Insights into the Lottery Ticket Hypothesis and Iterative Magnitude Pruning

@@ IV Results and Findings
$L$ Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$
1 $W^{Pr{(min\_(0))}}_{(1)}$ 1416.091 409.242
$W_{(1)}^{(min\_(1))}$ 1590.829 407.417
2 $W^{Pr{(min\_(1))}}_{(2)}$ 1512.426 402.678
$W_{(2)}^{(min\_(2))}$ 1349.951 399.680
3 $W^{Pr{(min\_(2))}}_{(3)}$ 1401.285 400.855
$W_{(3)}^{(min\_(3))}$ 1194.164 386.948
4 $W^{Pr{(min\_(3))}}_{(4)}$ 1266.607 392.829
$W_{(4)}^{(min\_(4))}$ 1129.884 375.590
5 $W^{Pr{(min\_(4))}}_{(5)}$ 17779.826 417.984
$W_{(5)}^{(min\_(5))}$ 1094.078 372.753
6 $W^{Pr{(min\_(5))}}_{(6)}$ 1590.785 420.431
$W_{(6)}^{(min\_(6))}$ 1065.735 373.472
7 $W^{Pr{(min\_(6))}}_{(7)}$ 2368.243 446.758
$W_{(7)}^{(min\_(7))}$ 1065.881 375.401
8 $W^{Pr{(min\_(7))}}_{(8)}$ 1882.278 432.505
$W_{(8)}^{(min\_(8))}$ 1200.277 382.051
9 $W^{Pr{(min\_(8))}}_{(9)}$ 3753.991 488.004
$W_{(9)}^{(min\_(9))}$ 1334.666 394.166
10 $W^{Pr{(min\_(9))}}_{(10)}$ 5426.848 525.621
$W_{(10)}^{(min\_(10))}$ 1575.910 408.810
@@
Solution Average Radius of Basin
$W_{(10)}^{(min\_(10))}$ 0.232314
$W^{Pr{(min\_(9))}}_{(10)}$ 0.183428
@@
$L$ Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$
1 $W^{RPr{(min\_(1))}}_{(0)}$ 1742.268 418.034
$W_{(0)}^{(min\_(0))}$ 1575.277 415.42
2 $W^{RPr{(min\_(2))}}_{(1)}$ 1572.214 415.826
$W_{(1)}^{(min\_(1))}$ 1590.829 407.417
3 $W^{RPr{(min\_(3))}}_{(2)}$ 1496.211 403.674
$W_{(2)}^{(min\_(2))}$ 1349.951 399.68
4 $W^{RPr{(min\_(4))}}_{(3)}$ 1326.676 394.347
$W_{(3)}^{(min\_(3))}$ 1194.164 386.948
5 $W^{RPr{(min\_(5))}}_{(4)}$ 1291.157 391.093
$W_{(4)}^{(min\_(4))}$ 1129.884 375.59
6 $W^{RPr{(min\_(6))}}_{(5)}$ 1316.275 391.471
$W_{(5)}^{(min\_(5))}$ 1094.078 372.753
7 $W^{RPr{(min\_(7))}}_{(6)}$ 1418.483 394.086
$W_{(6)}^{(min\_(6))}$ 1065.735 373.472
8 $W^{RPr{(min\_(8))}}_{(7)}$ 1464.987 398.559
$W_{(7)}^{(min\_(7))}$ 1065.881 375.401
9 $W^{RPr{(min\_(9))}}_{(8)}$ 1619.784 411.997
$W_{(8)}^{(min\_(8))}$ 1200.277 382.051
10 $W^{RPr{(min\_(10))}}_{(9)}$ 1913.214 427.475
$W_{(9)}^{(min\_(9))}$ 1334.66 394.166
@@
Solution Average Radius of Basin
$W_{(9)}^{(min\_(9))}$ 0.091809
$W^{RPr{(min\_(10))}}_{(9)}$ 0.074714
@@ IV-D Result 4: Why does the initialization proposed by the lottery ticket hypothesis work well?
Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$
$W^{(RIPN)}$ 6271.873 497.099
$W^{(min\_(10))}_{(10)}$ 1575.910 408.810
@@ IV-E Result 5: Why do we need the Iterative Process, and why does one-shot pruning not work as well?
Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$
$W^{(one\_shot)}$ 2944.881 450.409
$W^{(min\_(10))}_{(10)}$ 1575.910 408.810
@@ IV-F Result 6: What happens when you prune the smaller weights?
Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$
$W^{(RPN\_2)}$ 5635.883 487.644
$W^{(one\_shot)}$ 2944.881 450.409
@@
Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$
$W^{(RPN\_1)}$ 2100.56 435.624
$W^{(min\_(10))}_{(10)}$ 1575.91 408.81
@@ V Conclusion and Scope for Further Work
$L$ Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$
1 $W^{Pr{(min\_(0))}}_{(1)}$ 70.099 212.779
$W_{(1)}^{(min\_(1))}$ 50.77 203.681
6 $W^{Pr{(min\_(5))}}_{(6)}$ 86.27 198.106
$W_{(6)}^{(min\_(6))}$ 37.08 177.233
11 $W^{Pr{(min\_(10))}}_{(11)}$ 1413.523 414.844
$W_{(11)}^{(min\_(11))}$ 1326.092 352.86
@@
$L$ Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$
1 $W^{RPr{(min\_(1))}}_{(0)}$ 43.777 219.996
$W_{(0)}^{(min\_(0))}$ 67.929 220.915
6 $W^{RPr{(min\_(6))}}_{(5)}$ 46.733 201.835
$W_{(5)}^{(min\_(5))}$ 77.396 189.044
12 $W^{RPr{(min\_(12))}}_{(11)}$ 4379.201 426.294
$W_{(11)}^{(min\_(11))}$ 1326.092 352.86
@@
Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$
$W^{(RPN)}$ 51190.094 451.384
$W^{(min\_(12))}_{(12)}$ 3992.4783 389.957